OpenStudy (anonymous):

find the product: (x-1)(x-2)

OpenStudy (texaschic101):

are you familiar with the FOIL method ?

OpenStudy (anonymous):

I've heard of it

OpenStudy (texaschic101):

FOIL stand for first, outside, inside, last (x - 1)(x - 2) First -- multiply first terms in each set....x * x = x^2 Outside -- multiply outside terms in each set...x * -2 = -2x inside -- multiply inside terms...-1 * x = -x Last -- multiply last terms...-1 * -2 = 2 now put it all together... x^2 - 2x - x + 2 -- combine like terms x^2 - 3x + 2 any questions ?

OpenStudy (anonymous):

no thanks

OpenStudy (anonymous):

thankyou